# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v1

@@ 4.2 Idea generation through user interviews
The guided interviews consist of open-ended questions to learn about the stories of the interviewees’ daily routines. Interviewees in a focus group have expert knowledge of the logging system and represent most of the roles interacting with it, thus, ensuring different perspectives of the contextual inquiries. The focus group consists of seven persons with multiple roles in production: a plant coordinator, three shift leaders, a laboratory manager, a production assistant, and a second plant manager.
@@
    How do you use the logging system in your daily tasks? On which occasions? What information is of the most interest to you there?
@@
We summarize the interviews as follows: (1) as a timeline of a shift that generalizes three stories of the shift leaders, (2) as an overview of answers to the last two questions from all interviewees. These forms of summarization follow two different goals: first, to find the patterns in the user stories and identify which challenges or problems the interviewees mention at each step that we can turn into ideas for an NLP application. Second, to have an overview of the diverse opinions that can enable a broader picture for generating ideas.
@@
Figure 5: A generalization of a shift from a shift leader resulting from our interviews. The context inquiries identified three phases of a shift and the main tasks and events that may occur during the day. For the steps in all phases, we provide a short description and outline challenges, which we can turn into opportunities for an NLP application.
@@
Figure 6: An overview of the user replies for the questions from two rounds of interviews: to help create an idea for an NLP prototype and to evaluate a concept of the proposed prototype. From the interviewees with different roles, we learned about the diversity of use cases of the logging system. Although the interviewees described a few cases of data incompleteness, they agreed that using the information from the logged records as shared knowledge can support them on a daily basis.
@@
Figure 5 generalizes an average shift of a shift leader. We identify three phases of a shift, i.e., beginning, middle, and end of the shift, that determine which tasks a shift leader needs to perform. For the steps in the phases, we indicate a motivation for a step and which mentioned aspects we could turn into opportunities. For example, at the beginning of a shift, a shift leader reads through the logs of the past 24 hours to collect an overview of what happened over the time when they were not on duty. We learn, that if there is not enough information logged, then they talk to their colleagues, and the information from the oral communication remains unrecorded. Therefore, we mark this requirement for additional communication as an opportunity to improve data completeness in the logging system. We identified multiple cases throughout a shift which indicates that more data could be logged to assist in more informed daily operations. Additionally, we learned that when it comes to solving incidents, shift leaders rarely use the system to research a solution. The most common way to find information of interest is to read through the recent records or use a string-match search to find specific keywords, e.g., product names or order IDs. So we marked this aspect as an opportunity to facilitate knowledge sharing over the logging system.
@@
Figure 6 overviews how the interviewees use the logging system and their feedback about the data quality. On the one hand, the interviewees indicate a diverse spectrum of information they use from the system, e.g., acquire more context to the information from other systems running in the production, stay informed about deviation cases, read about results of specific experiments, or learn only about critical events over the past 30 days. On the other hand, the interviewees agree on the problems in the data quality and describe it as semi-complete. For example, the interviewees revealed that they often require oral communication to clarify unclear logged information. Additionally, they may need to refer to external sources of information to understand the full picture (e.g., emails, work chats, or attached files). Moreover, reporting style is person-dependent and sometimes lacks a reporting standard for specific types of events. The diverse writing styles may also depend on the type of reported event, i.e., reporting about some types of events required a higher level of detail, and reporting about others required only short messages due to using other systems for these events.
@@
The interviews confirmed our hypotheses about the high variation of the data quality and the rather short lifespan of the logged information. We learned from the interviewees that some records contained very detailed information, and some were rather scarce. We also learned that depending on the criticality of the events, interest in the records varies on average between 24h to 30 days. If the interviewees need to find any information, they apply available filters and read through a filtered list of logs. Occasionally, they use a string-match search to find records that contain very specific keywords that they have previously seen in the text.
@@
Our goal for the NLP application is to improve the exchange, i.e., documentation and retrieval of crucial information, which will ultimately lead to higher efficiency and shorter downtime of the plant. From the interviews, we learned about the common problems of data completeness, varying text quality, and the short life span of the records. To align with this goal and address the identified problems, we propose a concept of turning the logging system of daily operations into a source of shared domain knowledge (Figure 7). As of the current state, only the most recent logged information has value for the users. We assume that an NLP-driven domain knowledge processing system will enable leveraging the collected knowledge over the long run. Ultimately, users should not only generate information but also improve their domain expertise based on historical data.
Figure 7: The historic data of the plant operation contains a lot of know-how information on daily operations, but the use of it is quite limited due to the inefficient search. By implementing efficient semantic search and allowing users to find more complex records, the project aims at motivating the logging of more detailed descriptions of the daily operations, thus, improving the input data completeness and quality.
As a first stage, we propose a semantic search to enable efficient and effective access to the full history of daily operations. Our hypothesis is an efficient search engine for records should motivate users to log more complete and structured information. If the unstructured records are retrieved effectively, the new records with clear and complete information have more confidence to be found later on demand. Figure 7 depicts a goal of improving the data completeness in the system by logging more information, e.g., from personal oral and written communication. If the key points of the discussions can be later retrieved on demand, the users get more motivated to transfer this information into the logging system. Representation of the domain language is a core challenge for this domain NLP application. For the prototype phase, we require a solution that is feasible to design, implement, and test within a short period but also efficient and robust to perform user studies for the prototype evaluation. Section 5 reports more detail on this consideration.
4.3 Moderated concept evaluation
To evaluate our concept, we perform Wizard of Oz prototyping [10, 13], i.e., develop a “click-dummy” prototype, which mimics the functionality of the front-end of a semantic search engine. We adhere to the UX approach of “fail fast” and collect feedback on the proposed concept in the earliest possible stages of development.
@@ 4.3 Moderated concept evaluation
Figure 8: Based on the opportunities learned from the user interviews, we created a “click-dummy” prototype for a moderated concept evaluation through user interviews (Wizard of Oz prototyping). We demonstrated this prototype to the interviewees, asked them to interact with them, and ask if and how they would use it.
We conduct a second round of interviews where we present the click-dummy prototype and evaluate our concept via guided interviews and moderated usability sessions. In the one-to-one interviews, we aimed to learn if providing efficient access to historical data would allow more informed and effective daily operations:
@@
    Would an effective search engine assist you on a daily basis? If yes, how and when? Which queries would you devise?
@@
All interviewees agreed that efficient search would enable them to access the previously logged knowledge, e.g., solutions to problems. Moreover, they highlighted that a search application would provide an additional level of assistance at a night shift and could assist in planning product launching. The shared ideas for the queries helped us to create lists of queries for the later prototype evaluation (see Section 6). Figure 6, the last column, overviews the interviewees’ replies on how they would use a system of shared domain knowledge. For example, they highlighted a need to access solutions to previously occurring problems. Moreover, they expect to retrieve information on a level of deep understanding, e.g., for “quality state of product_A on Monday” they would expect to see information about the batch level on that day. Additionally, they would explore information in an aggregating way, e.g., assemble problems that occurred to a specific product.
5 Prototype Development
We develop a prototype of a semantic search engine for documents with uneven text structure and domain-specific language. Domain language includes professional terms, abbreviations, shortenings, chemical formulas, various IDs, and inventory codes of the plant machinery and equipment.
5.1 Pipeline
The search engine consists of two processes: (1) indexing of text records from a SQL database, (2) retrieving relevant records to a given query [^2] We use the terms “documents” to refer to the records of the logging system, i.e., short documents.. In the first prototype, we address a problem of the domain-specific language by using context expansion (i.e., adding definitions of some domain terms to the documents) and encoding documents with fastText (i.e., distributed word representation capable of handling out-of-vocabulary domain words).
@@ 5.1 Pipeline
Indexing starts with accessing a SQL database and reading values from string columns (Figure 9). We concatenate all text attributes of each record into one text document. Preprocessing performs so-called context expansion to add semantic descriptors to some domain terminology (Section 5.2). We transform documents into multiple vector spaces: (1) vectorization into boolean and term frequency-inverse document frequency (TF-IDF) representations that require full-text collection, (2) per-document vectorization into semantic representation, which employs word embedding service to represent words into fastText word vectors and computes TF-IDF weighted-mean word vectors to represent documents. Finally, we save boolean and semantic document representations into a search index. Document retrieval uses the same word embedding service and represents a query as a mean of word vectors (Figure 10). Candidate retrieval, first, retrieves documents that contain terms to be mandatory present in the documents (exact match), second, filters the documents by semantic similarity (Section 5.4). Lastly, we rank the top M most semantically similar documents by combined semantic similarity on the document and term levels.
5.2 Preprocessing
@@
Preprocessing focuses on (1) providing additional context for some domain-specific terms, i.e., dictionary-based context expansion, and (2) removing tokens that are irrelevant for semantic search based on the static word embeddings.
@@
To enable context expansion, we create a dictionary about functional locations, i.e., production machinery and equipment installed in a plant. The dictionary contains information about long IDs, their short equivalents, and descriptions. We create this dictionary based on a graph structure of the machinery from the logging system  [^3] We use the only available knowledge source from the logging system. Creating a dictionary or taxonomy for all domain terms (e.g., products) is impossible due to the tedious manual work of creating and maintaining such dictionaries. Many domain terms are described in a large variety of external systems that differ from plant to plant, so custom integration and updating are unrealistic..
We use the domain dictionary to extract descriptions of a mentioned machinery, e.g., in Figure 2, a description of “R105.12” is “reactor.” The dictionary serves to “translate” both long and short IDs of the machinery. Long IDs are used as attributes of the records, which the users attach when creating a record. The attributes indicate reporting about these specific functional locations. Moreover, short versions of IDs are used as in-text mentions of the same or additional functional locations. These in-text mentions typically look like abbreviations of the long codes and include both letters and digits.
@@
Descriptions of long and short IDs are added to the text before the occurrence of the ID. For the long IDs, we extract a description and a short ID of this piece of machinery and append them to the text at the very beginning, i.e., a title of a record to index. This approach is similar to automatic query expansion when query formulation is improved with additional sources of information [15]. Such context expansion enables retrieving records that either contain specific codes of functional locations, e.g., “R105.12”, or their descriptions, e.g., “reactor”.
The last step of preprocessing normalizes German text before encoding with static word embeddings. First, we perform tokenization, lemmatization, and POS-tagging with spaCy for German language [^4] https://spacy.io/models. By using the assigned POS tags, we remove tokens that belong to punctuation, articles, prepositions, and coordinating conjunctions. Second, we use lemmas to bring different word forms to a neutral form. Lastly, we did not lowercase text and used it as-is to enable a more accurate vector representation of nouns (the German language keeps nouns always capitalized).
5.3 Representation of the domain-specific language
Models for passage retrieval identify relevant paragraphs or short documents given questions or short keyword-based queries. One of the most common architectures for semantic document representation is sentence transformers [42]. A downside of pretrained models for passage retrieval based on sentence transformers is high sensitivity to the domain shift [49]. Domain adaption of transformer models is a set of techniques to make pretrained models learn peculiarities of domain language [26]. Domain-adaption of the transformer models for passage retrieval typically requires large datasets [46]. Some techniques utilize a combination of large labeled datasets of the general domain and small unlabeled domain datasets but require an annotated dataset for evaluation [48]. For the prototype development, we follow a UX principle “fail fast, fail often” and start with simpler techniques that require less time and resources.
@@ 5.3 Representation of the domain-specific language
Our semantic search is based on fastText word vectors that use a skip-gram model for distributed representation on the char ngram level [36]. Most of the word embedding models, such as word2vec and GloVe, contain only vectors for the words that were in a training corpus [35, 39]. Out-of-vocabulary (OOV) words pose a big problem when word embeddings are applied to specific domains, e.g., biology or chemistry. Some models propose training word embeddings from scratch on a domain-specific corpus (cf. [20]) or using subword information to learn the representation of the words’ parts to be able to represent later OOV words of this domain (cf. [53]).
The ngram word representation of fastText effectively creates word vectors for the out-of-vocabulary (OOV) words, i.e., words that typically occur in compound-based languages such as German and in domain-specific texts [8, 25] (e.g., in biology [22]). Consequently, ngram-level word representation enables vectorizing word shortening and makes them semantically similar to the original words. For example, the vocabulary of a trained fastText model does not contain the word “Temperaturschwankungen” (i.e., “temperature fluctuations”), but a fastText model can still represent this word so that it is semantically close to the words “Temperatur” and “Schwankungen.” Moreover, a shortening such as “Temp.Schwank.” will be vectorized and be semantically similar to “Temperaturschwankungen.” Such property of ngram-level representation is crucial to vectorize text documents with a writing style of hand-written notes. We did not consider using ELMo, i.e., contextualized word embeddings [40], because (1) word vector representation with ELMo depends on the word’s context, (2) word representation may suffer from incomplete grammar or syntax of the domain short texts in our dataset. Therefore, we used simpler static word embeddings, which can represent the writing style of hand-written notes more robustly and do not require word sense disambiguation within the same narrow domain.
We use the TF-IDF-weighted average of the word vectors to highlight the impact of the semantically most important tokens. Weighting word embeddings with IDF has successfully been applied for IR tasks [17, 23, 44, 29, 6, 2].
5.4 Retrieval and ranking
@@
Document retrieval uses a boolean search model to find exactly matching terms and reduces a search space. If a query contains no terms for the exact match, all documents in a collection are considered selected documents. If at least one term is found in the indexed vocabulary, we retrieve documents that contain these found terms.
From the candidate documents, we retrieve documents that are the most semantically related to a query. Candidate retrieval computes document similarity as cosine similarity between a vector representation of a query and vector representations of all candidate documents. A linear search is not very time-efficient, but in the current prototype phase, higher accuracy of the approach is more important than execution time. Search in the domain of daily operation is similar to Web search, i.e., only a few most relevant records are important for the users. Therefore, we retain the top $K=200$ most similar documents and proceed with ranking.
@@
Our ranking is based on a harmonic mean of document and term similarities. Term similarity prioritizes documents containing the query’s terms as-is and ranks higher the documents containing synonyms to the query’s terms. Term similarity is computed as a mean of the maximal similarities of the query’s terms to the document’s terms (Figure 11):
@@
where $QT$ is query terms, $|QT|$ is a number of query terms, $qt$ is a vector representation of a query term, $D$ is a document, and $dt$ is a document term. A harmonic mean balances general semantic relevance on the document level and precise similarity on a term level that allows synonyms or shortenings of the terms to gain a high rank. If an exception situation occurs that a query contains only terms for exact match, then the documents are ranked by their timestamp.
5.5 User interface
The user interface of our prototype resembles the one proposed for the concept evaluation (see Section 4.3). Figure 12 depicts an interface of the prototype that contains the text of the record, enables sorting by relevance and timestamp, and contains functionality to collect user feedback for the prototype evaluation. Additionally, the prototype provides a detailed view by clicking on a record of interest. For the first version of the prototype, we excluded the functionality of highlighting and visualization of the relevance score to prevent judgment bias during the user-based evaluation of the prototype accuracy (see Section 6.1).

### v4

@@ 4.1.2 Idea generation through user interviews
The guided interviews consisted of open-ended questions to learn about the stories of the interviewees’ daily routines. Interviewees in a focus group have expert knowledge of the logging system and represent most of the roles interacting with it, thus ensuring different perspectives of the contextual inquiries. The focus group consisted of seven persons with multiple roles in production: a plant coordinator, three shift leaders, a laboratory manager, a production assistant, and a second plant manager.
@@
    How do you use the logging system in your daily tasks? On which occasions? What information would you be interested in there?
@@
We summarized the interviews as follows: (1) a timeline of a shift that generalizes three stories of the shift leaders, (2) an overview of answers to the last two questions from all interviewees. These forms of summarization followed two different goals. First, to find the patterns in the user stories and identify which challenges or problems the interviewees mention at each step that we can turn into ideas for an NLP application. Second, to get an overview of the diverse opinions that can enable a broader picture for generating ideas.
@@
Figure 5: A generalization of a shift from a shift leader resulting from our interviews. The context inquiries identified three phases of a shift and the main tasks and events that may occur during the day. We provide a short description and outline challenges for the steps in all phases, which we can turn into opportunities for an NLP application.
@@
Figure 6: An overview of the user replies for the questions from two rounds of interviews: to help create an idea for an NLP prototype and evaluate a proposed prototype concept. From the interviewees with different roles, we learned about the diversity of use cases of the logging system. Although the interviewees described a few cases of data incompleteness, they agreed that using the information from the logged records as shared knowledge can support them daily.
@@
Figure 5 generalizes an average shift of a shift leader. We identify three phases of a shift, i.e., beginning, middle, and end of the shift that determine which tasks a shift leader needs to perform. For the steps in the phases, we indicated a motivation for a step and which mentioned aspects we could turn into opportunities. For example, at the beginning of a shift, a shift leader reads through the logs of the past 24 hours to collect an overview of what happened when they were not on duty. We learned that if not enough information was logged, they talked to their colleagues, and the information from the oral communication remained unrecorded. We noted that recording the additional information from communication was an opportunity to improve data completeness in the logging system. We identified multiple cases throughout a shift, which indicated that more data could be logged to assist in more informed daily operations. Additionally, we learned that shift leaders rarely use the system to research a solution for solving incidents. The most common way to find information of interest is to read through the recent records or use a string-match search to find specific keywords, e.g., product names or order IDs. This aspect would be an opportunity to facilitate knowledge sharing in the logging system.
@@
Figure 6 overviews how the interviewees use the logging system and their feedback about the data quality. On the one hand, the interviewees indicated a diverse spectrum of information they use from the system, e.g., acquiring more context to the information from other systems running in the production, staying informed about deviation cases, reading about results of specific experiments, or learning only about critical events over the past 30 days. On the other hand, the interviewees agreed on the problems in the data quality and described it as semi-complete. For example, the interviewees revealed that they often require oral communication to clarify unclear logged information. Additionally, they may have needed to refer to external sources of information to understand the full picture (e.g., emails, work chats, or attached files). Moreover, the reporting style was person-dependent and sometimes lacked a reporting standard for specific events. The diverse writing styles may have depended on the type of reported event, i.e., reporting about some events required a higher level of detail, whereas reporting about others required only short messages due to using other systems for these events.
@@
The interviews confirmed our hypotheses about the high data quality variation and the logged information’s short lifespan. We learned from the interviewees that some records contained very detailed information, and some were scarce. We also learned that depending on the criticality of the events, interest in the records varies on average between 24 hours and 30 days. If the interviewees needed to find any information, they applied available filters to the logged records and read through a filtered list. Occasionally, they use a string-match search to find records containing specific keywords they have previously seen in the text.
@@
Our goal for the NLP application was to improve the information exchange, i.e., logging and retrieving crucial information. We expected the effective and efficient information exchange to yield higher production efficiency and shorter plant downtime. From the interviews, we learned about the common problems of data completeness, varying text quality, and the short life span of the records. To align with this goal and address the identified problems, we proposed turning the logging system of daily operations into a source of shared domain knowledge (Figure 7). As of the current state, only the most recent logged information is important for the users. We assumed that an NLP-driven domain knowledge processing system would enable leveraging the collected knowledge over the long run. Ultimately, users should be encouraged to log more information and improve their domain expertise by learning from the logged experience.
Figure 7: The historical data of the plant operation contains a lot of know-how information on daily operations, but the use of it is quite limited due to the inefficient search. By implementing efficient semantic search and allowing users to find more complex records, the project aims at motivating the logging of more detailed descriptions of daily operations, thus improving the input data completeness and quality.
As a first stage, we proposed a semantic search to enable efficient and effective access to the full history of daily operations. Our hypothesis was an efficient search engine for records should motivate users to log more complete and structured information. If the unstructured records were retrieved effectively, the new records with clear and complete information would have more confidence to be found later on demand. Figure 7 depicts a goal of improving the data completeness in the system by logging more information, e.g., from personal oral and written communication. If the discussions’ key points can be later retrieved on demand, users get more motivated to transfer this information into the logging system. Representation of the domain language was a core challenge for this domain NLP application. For the prototype phase, we required a feasible solution to design, implement, and test within a short period but also efficient and robust to perform user studies for the prototype evaluation. Section 4.2 reports more detail on this consideration.
4.1.3 Moderated concept evaluation
To evaluate our concept, we performed Wizard of Oz prototyping [10, 13], i.e., developed a “click-dummy” prototype, which mimics the functionality of the front-end of a semantic search engine. We adhere to the UX principle of “fail fast” and collect feedback on the proposed concept in the earliest possible stages of development.
@@ 4.1.3 Moderated concept evaluation
Figure 8: Based on the opportunities learned from the user interviews, we created a “click-dummy” prototype for a moderated concept evaluation through user interviews (Wizard of Oz prototyping). We demonstrated this prototype to the interviewees and inquired if and how they would use it.
We conducted a second round of interviews where we presented the click-dummy prototype and evaluated our concept via guided interviews and moderated usability sessions. In the one-to-one interviews, we aimed to learn if providing efficient access to historical data would allow perform daily tasks in a more informed and effective manner:
@@
    Would an effective search engine assist you daily? If yes, how and when? Which queries would you devise?
@@
All interviewees agreed that efficient search would enable them to access the previously logged knowledge, e.g., solutions to problems. Moreover, they highlighted that a search application would provide additional assistance on a night shift and could assist in planning product launches. The shared ideas for the queries helped us to create lists of queries for the later prototype evaluation (see Section 4.3). Figure 6, the last column, overviews the interviewees’ replies on how they would use a system of shared domain knowledge. For example, they highlighted a need to access solutions to previously occurring problems. Moreover, they expected to retrieve information on a level of deep understanding. For example, for a query “quality state of product_A on Monday,” they would expect to see information about the batch level on that day. Additionally, they would explore information in an aggregating way, e.g., assemble problems that occurred to a specific product.
4.2 Prototyping
The exploration phase resulted in a decision to develop a semantic search engine prototype for documents with uneven text structure and domain-specific language. The prototype’s main challenge is handling the domain-specific language, including professional terms, abbreviations, shortenings, chemical formulas, various IDs, and plant machinery and equipment inventory codes. For the prototype development, we followed a UX principle “fail fast, fail often” [19] and implemented simpler approaches that require less time and resources.
4.2.1 Pipeline
The search engine consists of two processes: (1) indexing of text records from a SQL database, (2) retrieving relevant records to a given query [^2] We use the terms “documents” to refer to the records of the logging system, i.e., short documents.. In the first prototype, we address a problem of the domain-specific language by using context expansion (i.e., adding definitions of some domain terms to the documents) and encoding documents with fastText (i.e., vector word representation capable of handling out-of-vocabulary domain words).
@@ 4.2.1 Pipeline
Indexing starts with accessing an SQL database and reading values from string columns (Figure 9). We concatenate all text fields of each record into one text document. Preprocessing performs context expansion to add semantic descriptors to some domain terminology (Section 4.2.2). We transform documents into multiple vector spaces: (1) vectorization into boolean and term frequency-inverse document frequency (TF-IDF) representations that require full-text collection, (2) per-document vectorization into semantic representation, which employs word embedding service to represent words into fastText word vectors and computes TF-IDF weighted-mean word vectors to represent documents. Finally, we save boolean and semantic document representations into a search index. Document retrieval uses the same word embedding service and represents a query as a mean of word vectors (Figure 10). First, candidate retrieval retrieves documents that contain terms to be mandatory present in the documents (depending on if a query contains terms in quotes). Second, it filters the documents by semantic similarity (Section 4.2.4). Lastly, we rank the top M most semantically similar documents by combined semantic similarity on the document and term levels.
4.2.2 Preprocessing
@@
Preprocessing focuses on (1) providing additional context for some domain-specific terms, i.e., dictionary-based context expansion, and (2) normalizing text for semantic search that employs static word embeddings, e.g., lowercasing, stopword removal, and lemmatization.
@@
To enable context expansion, we create a dictionary about functional locations, i.e., production machinery and equipment installed in a plant. The dictionary contains information about long IDs, their short equivalents, and descriptions. We created this dictionary based on a graph structure of the machinery from the logging system.
@@
We use the domain dictionary to extract descriptions of a mentioned machinery, e.g., in Figure 2, a description of “R105.12” is “reactor.” The dictionary serves to “translate” machinery codes. Descriptions of the codes are added to the text before their occurrence. This approach is similar to automatic query expansion when query formulation is improved with additional information [15]. Such context expansion enables retrieving records by containing either specific codes of functional locations or their descriptions.
4.2.3 Representation of the domain-specific language
@@
Our semantic search is based on fastText word vectors, i.e., a skip-gram model for distributed representation on the char n-gram level [42, 7, 27]. Most of the word embedding models, such as word2vec [41] and GloVe [46], contain only vectors for the words that were in a training corpus. Out-of-vocabulary (OOV) words pose a big problem when word embeddings are applied to specific domains, e.g., biology or chemistry [23]. Some models propose training word embeddings from scratch on a domain-specific corpus (cf. [21]) or using subword information to learn the representation of the words’ parts to be able to represent later OOV words of this domain (cf. [60]). But we used a pre-trained fastText model because the vectorization on the n-gram level was sufficient to represent German compound terms of this domain [^3] For example, the vocabulary of a trained fastText model does not contain the word “Temperaturschwankungen” (i.e., “temperature fluctuations”). However, a fastText model can still represent this word so that it is semantically close to the words “Temperatur” and “Schwankungen.” Moreover, a shortening such as “Temp.Schwank.” will be vectorized and be semantically similar to “Temperaturschwankungen.” . We use the TF-IDF-weighted average of the word vectors to highlight the impact of the semantically most important tokens. Weighting word embeddings with IDF has successfully been applied for IR tasks [17, 24, 50, 33, 5, 1].
We did not use contextualized word embeddings, e.g., language models, because word representation may suffer from incomplete grammar or syntax of the domain short texts in our dataset. Therefore, we used simpler static word embeddings, which can more robustly represent the writing style of hand-written notes and do not require word sense disambiguation within the same narrow domain. Later prototype improvements will be based on the domain adaptation of the dense passage retrieval based on sentence transformers [48, 56, 28, 52, 55].
4.2.4 Retrieval and ranking
@@
From the candidate documents, we retrieve documents that are the most semantically related to a query. Candidate retrieval computes document similarity as cosine similarity between a vector representation of a query and vector representations of all candidate documents. A linear search is not time-efficient, but in the current prototype phase, higher accuracy of the approach is more important than execution time. Search in the domain of daily operation is similar to Web search, i.e., only a few of the most relevant records are important for the users. Therefore, we retain the top $K=200$ most similar documents and proceed with ranking.
@@
Our ranking is based on a harmonic mean of document and term similarities. Term similarity prioritizes documents containing the query’s terms as-is and ranks the documents containing synonyms higher than the query’s terms. Term similarity is computed as a mean of the maximal similarities of the query’s terms to the document’s terms (Figure 11):
@@
where $QT$ is query terms, $|QT|$ is a number of query terms, $qt$ is a vector representation of a query term, $D$ is a document, and $dt$ is a document term. A harmonic mean balances general semantic relevance on the document level and precise similarity on a term level that allows synonyms or shortenings of the terms to gain a high rank. If an exception situation occurs that a query contains only terms for an exact match, then the documents are ranked by their timestamp.
4.2.5 User interface
The user interface of our prototype resembles a click-dummy prototype proposed for the concept evaluation (see Section 4.1.3). Figure 12 depicts an interface of the prototype that contains the text of the record, enables sorting by relevance and timestamp, and contains functionality to collect user feedback for the prototype evaluation. Additionally, the prototype provides a detailed view by clicking on a record of interest. For the first version of the prototype, we excluded the functionality of highlighting and visualization of the relevance score to prevent judgment bias during the user-based evaluation of the prototype accuracy (see Section 4.3.1).